# Superfluid Band Theory for the Rod Phase in the Magnetized Inner Crust Matter: Entrainment, Spin-orbit, Spin-triplet Pairing

### v1

@@ 2.2 Band theory for inner crust matter
The central idea of band theory is encapsulated in Bloch’s theorem,
@@
where $\hat{h}_{q,\bm{k}}$ is obtained from $\hat{h}_{q}$ by the standard replacement of derivatives,
@@
In the present study, we focus on the two-dimensional crystalline configuration corresponding to the rod (“spaghetti”) phase in the neutron-star inner crust. The system is assumed to be uniform along the $z$ direction, and hence the cell-periodic functions $\tilde{\psi}_{i\bm{k}}(\bm{r})$ depend only on the transverse coordinates $(x,y)$. As a consequence, spatial derivatives with respect to $z$ acting on $\tilde{\psi}_{i\bm{k}}$ vanish, while the longitudinal Bloch wave number $k_{\parallel}$ associated with the $z$ direction is retained. In this formulation, the normalization is imposed on the full Bloch wave functions $\psi_{i\bm{k}}(\bm{r})$, rather than on the cell-periodic part $\tilde{\psi}_{i\bm{k}}(\bm{r})$. Specifically, the normalization is carried out over a supercell consisting of $N_{k_{x}}\times N_{k_{y}}$ unit cells in the $x$ and $y$ directions, respectively. Under this convention, the normalization condition for the cell-periodic function $\tilde{\psi}_{i\bm{k}}$ takes the form
$\int dx\,dy\,\tilde{\psi}_{i\bm{k}}(x,y)=aN_{k_{x}}bN_{k_{y}}.$ (19)
where $a$ and $b$ denote the lattice constants along the $x$ and $y$ directions. In the present calculations, we restrict ourselves to a square lattice, $a=b$. With this choice of normalization, local densities are defined as spatial averages over the supercell. Accordingly, explicit factors of $(aN_{k_{x}})(bN_{k_{y}})$ appear in the expressions for the densities, reflecting the normalization volume of the Bloch states. Derivatives acting on the Bloch wave functions are evaluated using the shifted operator (18). For example, the particle density, kinetic energy density, and current density are given by
@@
$n^{f}_{n}=\frac{1}{(aN_{k_{x}})(bN_{k_{y}})}\sum_{i\bm{k}}\frac{dk_{\parallel}}{2\pi}\theta(\mu_{n}-\epsilon^{(n)}_{i\bm{k}})\theta(\epsilon^{(n)}_{i\bm{k}}-U^{(\text{max})}_{n})$, (25)
@@
$\mathcal{K}^{\mu\nu}=\frac{1}{(aN_{k_{x}})(bN_{k_{y}})}\sum_{i\bm{k}}\frac{dk_{\parallel}}{2\pi}\frac{\partial^{2}\epsilon^{(n)}_{i\bm{k}}}{\partial k_{\mu}\partial k_{\nu}}\theta(\mu_{n}-\epsilon^{(n)}_{i\bm{k}})$. (27)
The curvature of the band dispersion reflects the mobility of quasiparticles in the crystal, and the conduction density therefore measures the fraction of dripped neutrons that can move effectively in the presence of band-structure effects. Although we previously proposed a time-dependent approach to extract the effective mass directly Yoshimura and Sekizawa (2024), we employ the above standard band-theoretical definition in the present work, because our primary interest here is the modification of single-particle band structures induced by magnetic-fields and spin-orbit coupling.
@@ 2.3 Hartree-Fock-Bogoliubov theory
In Hartree-Fock-Bogoliubov (HFB) theory Ring and Schuck (2004), quasiparticle creation and annihilation operators $\hat{\beta}^{\dagger}$ and $\hat{\beta}$ are defined through the Bogoliubov transformation of the particle operators $\hat{a}^{\dagger}$ and $\hat{a}$ as
@@
Owing to the fermionic anti-commutation relations, the anomalous density satisfies the antisymmetry condition
@@
Here, $\kappa_{\sigma\sigma^{\prime}}$ denotes the anomalous density matrix in spin space, while $\kappa(\bm{r})$ represents the pairing order parameter expressed as a $2\times 2$ spin matrix. The pairing order parameter can be decomposed into spin-singlet and spin-triplet components Sigrist and Ueda (1991); Sigrist (2005),
@@
where the trace is taken over spin indices. Assuming that only isovector ($T=1$) pairing channels ($nn$ and $pp$) are relevant, the allowed pairing modes are restricted to $s$-wave spin-singlet and $p$-wave spin-triplet components. The singlet anomalous density reduces to the familiar expression
$\kappa^{(s)}(\bm{r})=\frac{1}{2}\quantity(\kappa_{{\!\uparrow}{\!\downarrow}}(\bm{r})-\kappa_{{\!\downarrow}{\!\uparrow}}(\bm{r}))=\kappa_{{\!\uparrow}{\!\downarrow}}(\bm{r})$ (39)
@@
where $\tilde{J}^{M=0}_{t}$, $\tilde{J}^{M=1}_{t}$, and $\tilde{J}^{M=2}_{t}$ correspond to the scalar, vector, and traceless symmetric components of the pairing spin-current, respectively. Although these coupling constants are constrained as $\tilde{C}^{J}_{t}=3\tilde{C}^{J0}_{t}=2\tilde{C}^{J1}_{t}=\tilde{C}^{J2}_{t}$ in principle, we treat them as independent parameters in the present study to allow for general spin-triplet pairing interactions. The specific values of these constants used in this study will be provided in the later section. Under the spherical symmetric systems, the decomposed spin-current components $\tilde{J}^{M}_{t}$ are connected with the order parameters of the $^{3}P_{M}$ pairing, respectively. However, the present rod phase does not have such symmetry and it is not clear whether these pairing spin-current components and order parameters can be exactly related. This is why we call these contributions simply rank-0, rank-1 and rank-2 components, respectively. Nevertheless, we note that it is possible that they could be approximately connected even in the non-spherical rod phases.
@@
where $\hat{h}$ and $\Delta$ are $2\times 2$ matrices in spin space. The spin components of the pairing field $\Delta$ are also reduced to the spin-singlet and triplet parts, which is
@@
$(i\sigma_{y}\bm{\sigma})\cdot\bm{\Delta^{(t)}}$ $=-\frac{i}{2}\quantity[\gradient\cdot\quantity[\tilde{\mathsf{B}}\cdot(i\sigma_{y}\bm{\sigma})]+\quantity[\tilde{\mathsf{B}}\cdot(i\sigma_{y}\bm{\sigma})]\cdot\gradient].$ (47)
where the pairing mean field $\tilde{\mathsf{B}}$ is obtained from the functional derivative of the spin-triplet pairing energy,
@@
where $\tilde{J}^{M}_{t}$ corresponds to the $^{3}P_{M}$ components of the pairing spin-current. The scale factor $R$ is fixed to $R^{2}=1$ in the present study.
2.4 Extension to finite-magnetic field
The HF and HFB frameworks can be extended to systems under strong magnetic-fields by taking into account two distinct effects: (i) the modification of the nuclear single-particle Hamiltonian, and (ii) the Landau quantization of background electrons. The detailed formalism has been presented in our previous work Yoshimura and Sekizawa (2025), and here we briefly summarize the essential ingredients relevant to the present study.
The single-particle Hamiltonian in the presence of a magnetic field is written as
@@ 2.4 Extension to finite-magnetic field
, where $\hat{h}^{(0)}_{q,\sigma\sigma^{\prime}}$ denotes the Hamiltonian in the absence of magnetic-fields, and $\hat{h}^{(B)}_{q,\sigma\sigma^{\prime}}$ represents the magnetic contribution. The latter consists of orbital and spin couplings,
$\hat{h}^{(B)}_{q,\sigma\sigma^{\prime}}=-\quantity(\bm{l}\delta_{q,p}+g_{q}\frac{\bm{\sigma}}{2})\bm{\cdot}\tilde{\bm{B}}_{q}.$ (52)
where $\bm{l}$ is the orbital angular-momentum operator and $\bm{\sigma}$ denotes the Pauli matrices. The orbital term acts only on protons, while the second term describes the Zeeman coupling of the nucleon spin to the magnetic field. In the present work, we neglect the orbital coupling of protons, since our primary interest lies in magnetic effects on neutron superfluidity and pairing correlations, and the orbital contribution is expected to play a subdominant role. The nucleon $g$-factors are taken as
@@
Since we are concerned with extremely strong magnetic fields, it is convenient to introduce the scaled magnetic field
@@
In principle, the magnetic field is a three-dimensional vector. However, owing to the symmetry of the rod phase, the $x$ and $y$ directions are equivalent. Accordingly, we consider magnetic fields oriented either along the $x$ direction (perpendicular to the rods) or along the $z$ direction (parallel to the rods).
@@
The electron chemical potential $\mu_{e}$ is determined self-consistently by imposing both the $\beta$-equilibrium condition, $\mu_{n}=\mu_{p}+\mu_{e}$, and the charge-neutrality condition,
@@ 3.1 General Settings
In this study, we develop a dedicated numerical code to perform superfluid band calculations for two-dimensional crystalline structures. Assuming translational invariance along the $z$ direction, the spatial domain in the $x$ and $y$ directions is discretized on a uniform grid with $N_{x}=N_{y}=24$ points and grid spacings $\Delta x=\Delta y=1~\mathrm{fm}$. Accordingly, the size of the unit cell is fixed to $L_{x}=L_{y}=24~\mathrm{fm}$ throughout this work. In momentum space, the plane-wave wavenumbers are discretized into one hundred points with a spacing of $\Delta k=0.015~\mathrm{fm}^{-1}$ up to a cutoff $k_{\mathrm{max}}=1.5~\mathrm{fm}^{-1}$. For the Bloch wave vectors, the two-dimensional Brillouin zone is discretized using a uniform mesh of $N_{k_{x}}=N_{k_{y}}=32$ points. A square lattice configuration is assumed for the crystal structure. Single-particle wave functions are obtained by diagonalizing the HF or HFB Hamiltonian, and the calculations are iterated self-consistently until convergence of the total energy is achieved. Spatial derivatives and the Poisson equation are evaluated using the Fast Fourier Transform (FFT) algorithm. For density mixing at each iteration step, the modified Broyden method is adopted. We use the SLy4 Chabanat et al. (1997, 1998) parameter set for the Skyrme EDF. We employ the SLy4 parameter set Chabanat et al. (1997, 1998) for the Skyrme energy density functional. The baryon density is fixed at $n_{B}=0.04~\mathrm{fm}^{-3}$, and the $\beta$-equilibrium condition described in the previous section is imposed self-consistently at each iteration step. Regarding the coupling constants of the pairing spin-current terms, there is presently no consensus on their appropriate values. In the present study, for simplicity and to isolate qualitative effects, we adopt identical coupling constants for all tensor components, $\tilde{C}_{t}^{J0}=\tilde{C}_{t}^{J1}=\tilde{C}_{t}^{J2}=60\,{\mathrm{MeV}}$.
@@ 3.2 Structure of Study
This study addresses several closely related issues: namely, (i) the entrainment effect under strong magnetic-fields, (ii) the combined influence of spin–orbit coupling and magnetic-fields on spin polarization and pairing phases, and (iii) the role of spin-triplet pairing interactions in the presence of magnetic-fields. Accordingly, different forms of the Hamiltonian are employed in different parts of the analysis. In the first place, the entrainment effect is investigated within the Hartree-Fock framework, using the Hamiltonian given in Eq. (14), where pairing correlations are neglected. The neutron effective mass is defined through Eqs. (24)–(27), and its dependence on the magnetic-field strength is systematically examined. In the second place, superfluid effects are included to study the response of spin-polarization and spin-singlet pairing to external magnetic-fields. In this part, two calculation settings are considered: one including the spin-orbit interaction and the other neglecting it, in order to clarify its role. In the third place, the emergence of spin-triplet superfluidity is investigated by performing calculations both with and without the spin-triplet pairing energy density functional. For each configuration, calculations are carried out over a range of magnetic-field strengths to explore how the condensation of spin-triplet components depends on external conditions.
These calculations require substantial computational resources. For the Hartree-Fock calculations, CPU parallelization is implemented over the Bloch wave vectors and plane-wave momenta $\bm{k}$. In the Hartree-Fock-Bogoliubov calculations, the computational cost is dominated by the diagonalization of large $4N\times 4N$ Hamiltonian matrices. To accelerate this procedure, we employ GPU parallelization using the CUDA framework for Hermitian matrix operations. To reduce the overall computational cost, a part of the calculations is first performed without explicitly including the Bloch wave vectors $k_{x}$ and $k_{y}$. The impact of fully incorporating band-structure effects is examined and discussed in the later sections of this paper.
@@ 4 RESULTS and DISCUSSION
In this section, we present and discuss the numerical results, organized according to their physical implications In the first place, we investigate the impact of the magnetic field on the neutron effective mass within the Hartree-Fock framework, excluding pairing correlations. In the second place, we include superfluidity to the response of the spin-singlet pairing and spin-polarization to magnetic-fields. In the third place, we analyze how spin-triplet condensations are modified by the presense of the spin-triplet pairing interactions and magnetic-fields.
4.1 On Entrainment

### v2

@@ 2.2 Band theory for inner crust matter
The central idea of band theory is encapsulated in the Bloch’s theorem,
@@
where $\hat{h}_{q,\bm{k}}$ can be obtained from $\hat{h}_{q}$ by replacing derivatives as
@@
In the present study, we focus on the two-dimensional crystalline configuration corresponding to the rod (or “spaghetti”) phase in the neutron-star inner crust. The system is assumed to be uniform along the $z$ direction, and hence the cell-periodic functions $\tilde{\psi}_{i\bm{k}}(\bm{r})$ depend only on the transverse coordinates $(x,y)$. As a consequence, spatial derivatives with respect to $z$ acting on $\tilde{\psi}_{i\bm{k}}$ vanish, while the longitudinal Bloch wave number $k_{\parallel}$ associated with the $z$ direction is retained. In this formulation, the normalization is imposed on the full Bloch wave functions $\psi_{i\bm{k}}(\bm{r})$, rather than on the cell-periodic part $\tilde{\psi}_{i\bm{k}}(\bm{r})$. Specifically, the normalization is carried out over a supercell consisting of $N_{k_{x}}\times N_{k_{y}}$ unit cells in the $x$ and $y$ directions, respectively. Under this convention, the normalization condition for the cell-periodic function $\tilde{\psi}_{i\bm{k}}$ takes the form
[…]
where $a$ and $b$ denote the lattice constants along the $x$ and $y$ directions, respectively. In the present calculations, we restrict ourselves to a square lattice, i.e., $a=b$. With this choice of normalization, local densities are defined as spatial averages over the supercell. Accordingly, explicit factors of $(aN_{k_{x}})(bN_{k_{y}})$ appear in the expressions for the densities, reflecting the normalization volume of the Bloch states. Derivatives acting on the Bloch wave functions are evaluated using the shifted operator (18). For example, the particle density, kinetic energy density, and current density are given by
@@
[…]
@@
[…]
The curvature of the band dispersion reflects the mobility of particles in the crystal, and the conduction density therefore measures the fraction of dripped neutrons that can move effectively in the presence of band-structure effects. Although we previously proposed a time-dependent approach to extract the effective mass directly Yoshimura and Sekizawa (2024), we employ the above standard band-theoretical definition in the present work, because our primary interest here is the modification of single-particle band structures induced by magnetic-fields and spin-orbit coupling.
@@ 2.3 Hartree-Fock-Bogoliubov theory
In the HFB theory Ring and Schuck (2004), quasiparticle creation and annihilation operators $\hat{\beta}^{\dagger}$ and $\hat{\beta}$ are defined through the Bogoliubov transformation of the particle operators $\hat{a}^{\dagger}$ and $\hat{a}$ as
@@
Owing to the fermionic anti-commutation relations, the anomalous density satisfies the antisymmetric condition
@@
Here, $\kappa_{\sigma\sigma^{\prime}}$ denotes the anomalous density matrix in the spin space, while $\kappa(\bm{r})$ represents the pairing order parameter expressed as a $2\times 2$ spin matrix. The pairing order parameter can be decomposed into spin-singlet and spin-triplet components Sigrist and Ueda (1991); Sigrist (2005),
@@
where the trace is taken over spin indices. Assuming that only isovector ($T=1$) pairing channels ($nn$ and $pp$) are relevant, the allowed pairing modes are restricted to $s$-wave spin-singlet and $p$-wave spin-triplet components. The singlet anomalous density reduces to the familiar expression,
$\kappa^{(s)}(\bm{r})=\frac{1}{2}\quantity(\kappa_{{\!\uparrow}{\!\downarrow}}(\bm{r})-\kappa_{{\!\downarrow}{\!\uparrow}}(\bm{r}))=\kappa_{{\!\uparrow}{\!\downarrow}}(\bm{r}).$ (39)
@@
where $\tilde{J}^{M=0}_{t}$, $\tilde{J}^{M=1}_{t}$, and $\tilde{J}^{M=2}_{t}$ correspond to the scalar, vector, and traceless symmetric components of the pairing spin-current, respectively. Although these coupling constants are constrained as $\tilde{C}^{J}_{t}=3\tilde{C}^{J0}_{t}=2\tilde{C}^{J1}_{t}=\tilde{C}^{J2}_{t}$, we treat them as independent parameters in the present study to allow for general spin-triplet pairing interactions. The specific values of these constants used in this study will be provided in the later section. Under the spherical symmetric systems, the decomposed spin-current components $\tilde{J}^{M}_{t}$ are connected with the order parameters of the $^{3}\text{P}_{M}$ pairing. However, the present rod phase does not have such symmetry and it is not clear whether these pairing spin-current components and order parameters can be exactly related. This is why we call these contributions simply rank-0, rank-1, and rank-2 components, respectively. Nevertheless, we note that it is possible that they could be approximately connected even in the non-spherical rod phases.
@@
where $\hat{h}$ and $\Delta$ are $2\times 2$ matrices in the spin space, as well as wave functions $u$ and $v$ are spin vectors. The spin components of the pairing field $\Delta$ are also reduced to the spin-singlet and triplet parts, which is
@@
$(i\sigma_{y}\bm{\sigma})\cdot\bm{\Delta^{(t)}}$ $=-\frac{i}{2}\quantity[\gradient\cdot\quantity[\tilde{\mathsf{B}}\cdot(i\sigma_{y}\bm{\sigma})]+\quantity[\tilde{\mathsf{B}}\cdot(i\sigma_{y}\bm{\sigma})]\cdot\gradient],$ (47)
where the pairing mean field $\tilde{\mathsf{B}}$ is a spacial tensor and obtained from the functional derivative of the spin-triplet pairing energy,
@@
where $\tilde{J}^{M}_{t}$ corresponds to the $^{3}\text{P}_{M}$ components of the pairing spin-current. The scale factor $R$ is fixed to unity in the present study.
2.4 Extension for finite-magnetic field systems
The HF and HFB frameworks can be extended to systems under strong magnetic field by taking into account two distinct effects: (i) the modification of the nuclear single-particle Hamiltonian, and (ii) the Landau quantization of background electrons. The detailed formalism has been presented in our previous work Yoshimura and Sekizawa (2025), and here we briefly summarize the essential ingredients relevant to the present study.
The single-particle Hamiltonian in the presence of a magnetic-field is written as
@@ 2.4 Extension for finite-magnetic field systems
where $\hat{h}^{(0)}_{q,\sigma\sigma^{\prime}}$ denotes the Hamiltonian in the absence of magnetic field, while $\hat{h}^{(B)}_{q,\sigma\sigma^{\prime}}$ represents the magnetic contribution. The latter consists of orbital and spin couplings,
$\hat{h}^{(B)}_{q,\sigma\sigma^{\prime}}=-\quantity(\bm{l}\delta_{qp}+g_{q}\frac{\bm{\sigma}}{2})\bm{\cdot}\tilde{\bm{B}}_{q}.$ (52)
where $\bm{l}$ is the orbital angular momentum operator and $\bm{\sigma}$ denotes the Pauli matrices. The orbital term acts only on protons, while the second term describes the Zeeman coupling of the nucleon spin to the magnetic field. In the present work, we neglect the orbital coupling of protons, since our primary interest lies in magnetic effects on neutron superfluidity and pairing correlations, and the orbital contribution is expected to play a minor role. The nucleon $g$-factors are given as
@@
Since we are concerned with extremely strong magnetic fields, it is customary to introduce a scaled magnetic field as
@@
In principle, the magnetic field is a three-dimensional vector. However, owing to the symmetry of the rod phase, the $x$ and $y$ directions are equivalent. Accordingly, we consider magnetic fields oriented either along the $x$ direction (perpendicular to the rods) or along the $z$ direction (parallel to the rods), in the present study.
@@
The electron chemical potential $\mu_{e}$ is determined self-consistently by imposing both the $\beta$-equilibrium condition, $\mu_{n}=\mu_{p}+\mu_{e}$, and the charge neutrality condition,
@@ 3.1 General settings
In this study, we develop a dedicated numerical code to perform superfluid band calculations for two-dimensional crystalline structures. Assuming translational invariance along the $z$ direction, the spatial domain in the $x$ and $y$ directions is discretized on a uniform grid with $N_{x}=N_{y}=24$ points and a grid spacing of $\Delta x=\Delta y=1~\mathrm{fm}$. Accordingly, the size of the unit cell is fixed to $L_{x}=L_{y}=24~\mathrm{fm}$ throughout this work. In the momentum space, the plane-wave wavenumbers are discretized into one hundred points with a spacing of $\Delta k=0.015~\mathrm{fm}^{-1}$ up to a cutoff $k_{\mathrm{max}}=1.5~\mathrm{fm}^{-1}$. For the Bloch wave vectors, the two-dimensional Brillouin zone is discretized using a uniform mesh of $N_{k_{x}}=N_{k_{y}}=32$ points. A square lattice configuration is assumed for the crystalline structure. Single-particle wave functions are obtained by diagonalizing the HF or HFB Hamiltonian, and the calculations are iterated self-consistently until convergence of the total energy is achieved. Spatial derivatives and the Poisson equation are evaluated using the fast Fourier transform (FFT) algorithm. For the update of densities at each iteration step, the modified Broyden method is adopted. We use the SLy4 Chabanat et al. (1997, 1998) parameter set for the Skyrme EDF. The baryon density is fixed at $n_{B}=0.04~\mathrm{fm}^{-3}$, and the $\beta$-equilibrium condition described in the previous section is imposed self-consistently at each iteration step. Regarding the coupling constants of the pairing spin-current terms, there is presently no consensus on their appropriate values. In the present study, for simplicity and to isolate qualitative effects, we adopt identical coupling constants for all tensor components, i.e., $\tilde{C}_{t}^{J0}=\tilde{C}_{t}^{J1}=\tilde{C}_{t}^{J2}=60\,{\mathrm{MeV}}$.
@@ 3.2 Structure of study
This study addresses several closely related issues: namely, (i) the entrainment effect under strong magnetic-fields, (ii) the combined influence of spin-orbit coupling and magnetic fields on spin polarization and pairing phases, and (iii) the role of spin-triplet pairing interactions in the presence of magnetic-fields. Accordingly, different forms of the Hamiltonian are employed in different parts of the analysis. In the first place, the entrainment effect is investigated within the HF framework, using the Hamiltonian given in Eq. (14), where pairing correlations are neglected. The neutron effective mass is defined through Eqs. (24)–(27), and its dependence on the magnetic-field strength is systematically examined. In the second place, superfluid effects are included to study the response of spin-polarization and spin-singlet pairing to external magnetic-fields. In this part, two calculation settings are considered: one including the spin-orbit interaction and the other neglecting it, in order to clarify its role. In the third place, the emergence of spin-triplet superfluidity is investigated by performing calculations both with and without the spin-triplet pairing energy density functional. For each configuration, calculations are carried out over a range of magnetic-field strengths to explore how the condensation of spin-triplet components depends on external conditions.
These calculations require substantial computational resources. For the HF calculations, CPU parallelization is implemented over the Bloch wave vectors and plane-wave momenta $\bm{k}$. In the HFB calculations, the computational cost is dominated by the diagonalization of large $4N\times 4N$ Hamiltonian matrices. To accelerate this procedure, we employ GPU parallelization using the CUDA framework for Hermitian matrix operations. To reduce the overall computational cost, a part of the calculations is first performed without explicitly including the Bloch wave vectors $k_{x}$ and $k_{y}$. The impact of fully incorporating band-structure effects is examined and discussed in the later sections of this paper.
@@ 4 RESULTS and DISCUSSION
In this section, we present and discuss the numerical results, organized according to their physical implications. First, we investigate the impact of the magnetic field on the neutron effective mass within the HF framework, excluding pairing correlations. Secondly, we include superfluidity to the response of the spin-singlet pairing and spin-polarization to magnetic-fields. Finally, we analyze how spin-triplet condensations are modified by the presense of the spin-triplet pairing interactions and magnetic fields.
4.1 Entrainment effects